### v1

# A Risk Taxonomy for Evaluating AI-Powered Psychotherapy Agents

Abstract.
The proliferation of Large Language Models (LLMs) and Intelligent Virtual Agents acting as psychotherapists presents significant opportunities for expanding mental healthcare access. However, their deployment has also been linked to serious adverse outcomes, including user harm and suicide, facilitated by a lack of standardized evaluation methodologies capable of capturing the nuanced risks of therapeutic interaction. Current evaluation techniques lack the sensitivity to detect subtle changes in patient cognition and behavior during therapy sessions that may lead to subsequent decompensation. We introduce a novel risk taxonomy specifically designed for the systematic evaluation of conversational AI psychotherapists. Developed through an iterative process including review of the psychotherapy risk literature, qualitative interviews with clinical and legal experts, and alignment with established clinical criteria (e.g., DSM-5) and existing assessment tools (e.g., NEQ, UE-ATR), the taxonomy aims to provide a structured approach to identifying and assessing user/patient harms. We provide a high-level overview of this taxonomy, detailing its grounding, and discuss potential use cases. We discuss two use cases in detail: monitoring cognitive model-based risk factors during a counseling conversation to detect unsafe deviations, in both human-AI counseling sessions and in automated benchmarking of AI psychotherapists with simulated patients. The proposed taxonomy offers a foundational step towards establishing safer and more responsible innovation in the domain of AI-driven mental health support.
AI Psychotherapy, Risk Assessment, Intelligent Virtual Agents, LLM Safety, Mental Health Technology
journalyear: 2025conference: ; May 2025; booktitle: Preprintccs: Applied computing Consumer healthccs: Information systems Language modelsccs: Human-centered computing Empirical studies in HCI
1. Introduction
Artificial Intelligence (AI) systems, particularly those leveraging Large Language Models (LLMs), are increasingly being developed and deployed as tools for mental health support, including entertainment chatbots ”masquerading” as therapists (Jurblum and Selzer, 2025; Barry, 2025; Jin et al., 2025; Na et al., 2025). The American Psychological Association has recently warned federal regulators about AI chatbots programmed to reinforce, rather than challenge, a user’s thinking, potentially driving vulnerable users to harm themselves or others (Lawrence et al., 2024; Barry, 2025). Reports have emerged, including instances where interactions with AI chatbot ”therapists” have been linked to severe negative outcomes. Notably, one case involved a 14-year-old boy’s suicide after months of interaction with chatbots on Character.AI, including a chatbot that falsely claimed to have several years’ experience as a licensed therapist (Roose, 2024; Barry, 2025). In this instance, the chabot reportedly reinforced the user’s harmful ideations rather than challenging them (Barry, 2025). Such events underscore the profound ethical responsibilities and potential harms associated with deploying these technologies to vulnerable populations, often with concerns raised about them being ”dangerous and untested” or lacking rigorous, context-appropriate safety evaluations (Roose, 2024; Barry, 2025; Stade et al., 2024b).
A critical issue is the absence of standardized methods to evaluate these AI systems for the unique risks inherent in psychotherapeutic interactions. Existing LLM benchmarks often focus on general capabilities such as toxicity detection in isolated texts, failing to capture harms that may emerge from sustained, dynamic conversations (Chang et al., 2023; Zhang et al., 2025). Traditional clinical trials, while vital for efficacy, can be slow and ethically challenging for initial safety validation of unproven autonomous systems (Stade et al., 2024b; Heinz et al., 2025). This gap leaves users susceptible to unexpected adverse outcomes.
To address this, we propose a risk taxonomy tailored for evaluating AI psychotherapists. We present a high-level overview of this taxonomy, derived from an iterative process including a review of the literature on psychotherapy risks, qualitative interviews with clinical and legal experts, and alignment with established clinical criteria and existing assessment tools. We hypothesize that this taxonomy can provide a structured and systematic approach to identifying, categorizing, and ultimately mitigating risks in AI-driven psychotherapy applications. Our primary aim is to provide a framework that can be used to evaluate automated psychotherapy systems for potential negative impacts before and during their deployment, fostering safer innovation in digital mental health.
2. Related Work
2.1. Risks in Psychotherapy
Psychotherapy, while generally effective, can occasionally lead to negative consequences (Linden and Schermuly-Haupt, 2014; Boisvert and Faust, 2002). Estimates of unwanted effects vary from 3% to over 50%, depending on the particular definitions and assessment methods used (Linden, 2013; Mejía-Castrejón et al., 2024). A key challenge is the lack of a uniform conceptual framework and standardized tools for assessing adverse events (AEs) in psychotherapy (Mejía-Castrejón et al., 2024).
Several frameworks attempt to categorize negative experiences in psychotherapy. Linden and Schermuly-Haupt distinguish Unwanted Events (UEs) from Adverse Treatment Reactions (ATRs) caused by correctly applied treatment, and malpractice reactions from incorrectly applied treatment (Linden, 2013; Linden and Schermuly-Haupt, 2014). Mejía-Castrejón et al. (Mejía-Castrejón et al., 2024) adapted a medical model, classifying AEs by severity, relatedness, seriousness, and expectedness. Common AEs identified across studies and captured in assessment tools like the Negative Effects Questionnaire (NEQ) (Rozental et al., 2019) or the Unwanted Event to Adverse Treatment Reaction (UE-ATR) checklist (Linden, 2013) include negative emotions (e.g., anxiety, sadness), symptom worsening, the emergence of unpleasant memories, relationship strains, dependency on therapy, or reduced self-efficacy (Linden and Schermuly-Haupt, 2014; Yao et al., 2020). Factors influencing these include patient and therapist characteristics, and specific techniques (Linden and Schermuly-Haupt, 2014).
Assessing these effects is difficult; therapists may struggle to recognize negative outcomes, and establishing causality is complex (Linden, 2013; Linden and Schermuly-Haupt, 2014). While tools like the UE-ATR, NEQ, and the Inventory for Negative Effects of Psychotherapy (INEP) (Ladwig et al., 2014) exist, consensus on definitions and instruments remains elusive (Mejía-Castrejón et al., 2024). Reviews show variability and inadequate reporting of negative effects in randomized clinical trials (Jonsson et al., 2014; Klatte et al., 2023), underscoring the need for standardized assessment. Our taxonomy development aims to build upon these existing efforts by providing a structured framework specially adapted for evaluating AI-powered psychotherapy agents.
2.2. Safety Evaluation of LLMs in Healthcare
The application of LLMs in high-stakes domains like healthcare, including psychotherapy, demands rigorous safety assessment due to the potential for significant harm (Stade et al., 2024b; Lawrence et al., 2024; Steenstra et al., 2024). General LLM safety evaluations often focus on risks including bias, toxicity, or misinformation (Weidinger et al., 2021; Ji et al., 2023), using benchmarks such as ALERT (Tedeschi et al., 2024) or SafetyBench (Zhang et al., 2023). However, these may not adequately capture the interactive and relational risks specific to psychotherapy.
Safety frameworks are emerging for AI in mental health, including the READI framework (Stade et al., 2024a) and the BPES checklist, emphasizing ”Do No Harm” (Khan and Seto, 2023). Dialogue safety in mental health support is also a focus (Qiu et al., 2023). For instance, $\Psi$-ARENA offers an interactive platform for evaluating LLM-based psychological counselors (Zhu et al., 2025). Despite these advances, a standardized taxonomy to systematically identify and categorize the spectrum of potential risks specific to AI psychotherapeutic interactions, grounded in clinical realities and existing harm assessment principles, is still lacking. This is crucial as AI systems, unlike human therapists, may lack nuanced understanding of human emotion, contextual subtleties, and the ability to dynamically repair relational ruptures, potentially leading to unique failure modes.
3. Taxonomy Development
The development of our risk taxonomy followed an iterative design process. We first formulated a preliminary taxonomy based on an in-depth literature review of psychotherapy risk (e.g., (Linden, 2013; Rozental et al., 2018; Klatte et al., 2023)). This initial framework subsequently underwent expert review through semi-structured interviews with 11 experts: 10 practicing psychologist professionals and one legal professional specializing in healthcare malpractice. These interviews were designed to explore perceptions of risk in human therapy, identify potential failure modes for AI therapists, and gather direct feedback on the preliminary taxonomy. The interview study was approved by Northeastern University’s IRB.
The interviews were analyzed using thematic analysis (Clarke and Braun, 2017) to reveal six key themes.
3.1. Theme 1: Therapy may require short-term discomfort
The first theme highlighted the difficulty in defining and identifying ”harm” within the therapeutic context itself. Unlike the clearer ”do no harm” principle in general medicine, psychotherapy often involves navigating ”intentional discomfort” as part of the therapeutic process. Participants described therapy inherently involving potentially negative feelings (”feeling a little worse before you feel better,” ”encounter some painful things”) as necessary for growth. Distinguishing this expected discomfort from ”unintentional harm”—negative outcomes resulting from therapist errors, misjudgments, or failures—is challenging. Harm exists on a ”spectrum,” ranging from temporary setbacks (”bruises”) like feeling ”misunderstood” or ”judged,” to more significant consequences (”cuts”) such as a ”rupture” in the ”therapeutic alliance,” premature termination (”stop coming to therapy,” ”drop out early”), worsening symptoms (”exacerbate symptoms”), loss of ”trust in self,” developing a ”bad taste… about therapy,” or, in the most severe cases, increased ”suicide” risk. Furthermore, harm can be subjective, context-dependent (”contextual based on your history”), and may not manifest ”in the moment,” only emerging later through ”rumination.”
3.2. Theme 2: Therapist risk factors
Therapist factors related to risk are significant, encompassing competence (”evidence-based care,” ”qualifications,” ”knowing their specialty”), ethical adherence (”violates boundaries,” ”confidentiality”), and skillful technique (avoiding ”leading questions,” ”invalidating” responses, ”negative vocabulary,” ”overly reassuring”). Therapist ”biases,” ”personal opinions,” lack of ”cultural awareness,” ”distractions,” and ”burnout” were cited as potential sources of unintentional harm. Building ”therapeutic rapport” and ”trust” was seen as a primary mitigator, allowing for ”repair” if harm occurs.
3.3. Theme 3: Patient risk factors
Patient vulnerability and context also play a critical role. ”Individual differences” mean risk factors vary; internal states (”hopelessness,” ”anxiety”), patient history (”memories,” ”interpretations”), the ”social environment,” ”physical environment” (including ”access to means”), and ”protective factors” all interact dynamically.
3.4. Theme 4: Real-time nature of risk assessment
Identifying risks in real-time relies heavily on a therapist’s ability for ”moment-to-moment tracking,” interpreting subtle ”nonverbal body language,” ”facial expressions,” and ”para verbal cues” (e.g., ”intonation”), alongside verbal content and formal ”assessment tools” (e.g., ”Columbia Scale”(Posner et al., 2008), ”PHQ-9”(Kroenke et al., 2001)). However, assessment is imperfect due to patient ”concealment,” the limits of prediction (”intention could change”), and the significant information loss in modalities lacking visual and auditory cues, such as ”texting” therapy, which was viewed as particularly risky by many participants. Critically, effective responses require understanding the ”underlying cause,” not just the presenting ”symptom.”
3.5. Theme 5: Perceived shortcomings of AI therapists
When considering AI therapists, participants largely believed the fundamental risks in human therapy remain the same but are potentially ”higher” or exacerbated. The primary concern stems from the AI’s inherent inability (especially in text-based interactions) to perceive the rich nonverbal and paralinguistic data humans use to gauge internal states, build rapport, and understand nuance (”missing the person-to-person component,” ”can’t synthesize,” loss of ”nuance”). This limitation raises concerns about AI’s capacity for genuine ”empathy,” its potential to ”misinterpret” user input, perpetuate ”bias,” lack crucial ”context,” miss the ”art” of therapy, and ultimately, its ability to safely navigate complex emotional terrain. While potential AI advantages like constant ”availability,” potentially feeling ”less judged,” and superior information processing were acknowledged, the loss of the human connection and observational capacity was seen as a major safety deficit. Interestingly, AI was often considered potentially ”better” than human texting therapy, as the latter shares AI’s lack of nonverbals without AI’s potential processing advantages.
3.6. Theme 6: Refinement of risk taxonomy
The feedback on the proposed risk taxonomy indicated it was a ”good start” covering many ”big risks,” but highlighted needs for refinement. Key omissions noted were ”trauma” and ”abuse.” Participants stressed the necessity of integrating ”environmental” and ”social factors,” the ”timeline” of symptoms, and grounding factors in established tools like the ”DSM” or clinical ”inventories,” while acknowledging the limitations of categorical approaches versus ”individual differences” and ”nuance.” The taxonomy’s current structure was seen as potentially useful but insufficient alone for risk assessment without incorporating deeper context and validated measures.
These insights drove significant revisions: patient factors within the risk taxonomy were linked to the DSM-5, factor values are now compared to baselines to accommodate individual differences, and risk categorization was grounded in approaches similar to the NEQ and UE-ATR.
4. Overview of the Risk Taxonomy
The refined risk taxonomy, derived from literature (Linden, 2013; Rozental et al., 2018; Klatte et al., 2023), expert interviews, and DSM-5 criteria (Black and Grant, 2014), provides a structured framework for identifying potential negative outcomes in AI psychotherapy. The full taxonomy can be found in Appendix A. It has two main categories:
1. Immediate Risk: This category encompasses situations of imminent danger to the user or others, requiring urgent intervention. Examples include explicit suicidal intent with plan and means, threats of harm to others, or severe psychological decompensation such as acute psychosis. Identification relies primarily on dialogue-based triggers. Evaluation assesses the AI’s ability to recognize these triggers and respond according to predefined, ethically-informed action plans, such as attempting de-escalation, recommending emergency resources (e.g., 911, 988 Lifeline), and contacting/flagging a human supervisor immediately.
2. Potential Risk: This category refers to emerging states or interaction patterns suggesting heightened vulnerability to negative therapeutic outcomes, functional deterioration, or adverse impacts on the therapeutic process, often aligning with unwanted events in the NEQ or UE-ATR (Rozental et al., 2019; Linden, 2013). Key subcategories are: 2.1. Symptom Exacerbation, Emergence of New Symptoms, or Maladaptive Coping, which involves worsening of clinical symptoms (e.g., increased negative thoughts or feelings) or new maladaptive behaviors (e.g., Worry/Rumination Frequency); 2.2. Patient Destabilization or Feeling Overwhelmed, indicating the user is struggling with the therapeutic process (e.g., increased confusion, emotional lability); 2.3. Therapeutic Alliance Damage or Negative Relational Dynamics, covering negative shifts in the user-AI relationship including mistrust, feeling judged, or over-dependence; and 2.4. Reduced Engagement or Potential for Premature Termination, signaling decreased participation or commitment within the therapeutic process or current therpay session.
Identification of these potential risks involves tracking dynamic changes in internal cognitive, emotional, and behavioral factors, which are conceptualized as scales (e.g., from 1-low to 10-high for intensity or frequency, such as ”Hopelessness Level” or ”Collaboration/Engagement Level”). The up (↑) and down (↓) arrows accompanying these factors in the full taxonomy (see Appendix A) specify the direction of change from a user’s baseline state that may signify an increased potential risk. For example, an increase (↑) in ”Hopelessness Level” or a decrease (↓) in ”Collaboration/Engagement Level” would be flagged as potentially problematic.
Evaluation should assess harmful shifts in these factors from baseline values, allowing for the tracking of subtle deteriorations or emerging vulnerabilities. While empirically-grounded numeric thresholds for what constitutes a significant change is a goal for future work, the current approach suggests that any major deviation beyond simple, minor fluctuations around a baseline cognitive factor value (e.g., more than +/- 1 point on a 1-10 scale from the baseline), or other abnormal fluctuations, could indicate an increased potential risk requiring attention. Furthermore, as expert participants highlighted, these indicators should ideally be monitored over the course of a session or across multiple interactions to understand their trajectory and contextual significance, rather than relying solely on isolated measurements. For instance, a consistent decrease (↓) in the Motivation/Self-Efficacy Level factor over multiple sessions could strongly suggest an increased risk for Reduced Engagement / Potential for Premature Termination.
5. Expected Use Cases and Adaptability
The taxonomy standardizes safety evaluation criteria for AI-powered psychotherapy agents. There are several ways the taxonomy can be used to improve safety. We outline several use cases here.
5.1. Monitoring Real User Interactions
The risk taxonomy could be used to identify emergent risks in real-time, as a user interacts with an AI therapist, for example to provide early termination of a harmful session, or feedback to developers or regulatory bodies. Monitoring should be performed with ethical consent, and could be accomplished by tracking user risk factors from pre- to post-interaction to identify ”Potential Risk” shifts, while dialogue tools scan for ”Immediate Risk” triggers.
5.2. Evaluation with Simulated Patients
The risk taxonomy could also be used to ”monitor” interactions between an AI therapist and a simulated patient, such as (Steenstra et al., 2025), providing a powerful evaluation testbed that can be used to assess systems for safety and potential harm. During these simulated interactions, simulated patient agents’ dynamic internal states corresponding to ”Potential Risk” factors (e.g., ”Hopelessness Level) are updated. This allows systematic testing by comparing baseline to post-interaction values of internal patient factors and observing AI responses to diverse personas, a methodology previously validated in our development of SimPatient (Steenstra et al., 2025), an LLM-powered system where embodied simulated patients featured dynamic internal cognitive states (e.g., self-efficacy) that evolved during interactions.
5.3. Benchmarking and Comparative Analysis
The risk taxonomy provides a standardized framework essential for robust benchmarking and comparative analysis of different AI psychotherapist models or system versions. By systematically evaluating AI therapists against the defined risk categories and factors, consistent and comparable risk profiles can be generated for each system. This process can highlight the relative strengths and weaknesses of various AI approaches in terms of their propensity to elicit or mitigate specific types of risks. For example, one model might excel at de-escalating immediate suicidal ideation but perform poorly by increasing therapeutic alliance damage, while another might show the opposite pattern. Such comparative insights are invaluable for developers seeking to enhance model safety, for organizations making informed deployment decisions based on specific use-case risk tolerances, and for fostering a more transparent and accountable development ecosystem focused on minimizing harm.
5.4. Identifying Unexpected Outcomes
Finally, the risk taxonomy serves as a crucial tool for identifying unexpected or emergent adverse outcomes that may arise during automated therapy. This is achieved through systematic monitoring of user interactions and therapeutic outcomes, continually assessed against the taxonomy’s defined risk factors and categories. Such monitoring can reveal significant deviations from established baseline values—which ideally exhibit only safe fluctuations—or flag the emergence of previously unrecognized harmful patterns. When substantial deviations or novel harmful patterns are detected, a deeper analysis is warranted. This involves scrutinizing the conversational context surrounding these occurrences to understand the specific interaction dynamics or AI behaviors that may have precipitated them. Uncovering these root causes is invaluable, as it directly informs targeted safety improvements and iterative refinements of the AI model, ultimately leading to a more robust and safer therapeutic tool.
While our risk taxonomy provides a general framework, it can also be augmented for specific populations or problems. For instance, evaluating an AI psychotherapist for helping patients with substance use disorders could add ”Potential Risk” factors such as ”Substance Craving Intensity”, ”Ambivalence about changing substance use” ”Substance Use Frequency/Amount,” ”Level of Intoxication/Withdrawal Symptoms”, and relational dynamics like ”Dishonesty/Minimization regarding substance use.” This tailoring allows for more nuanced risk assessment. Future work includes validating the taxonomy against human expert judgments.
6. Conclusion
AI psychotherapists present significant opportunities and risks. Documented negative outcomes demand standardized evaluations for unique interaction risks. We developed the risk taxonomy, from an iterative process involving literature, expert insights, and DSM-5, and harm checklists. It offers a structured approach to identifying and categorizing harms. Its overview, grounding, and adaptable use cases (tracking risk factor changes from baseline to post-interaction) aim to equip stakeholders to ensure AI psychotherapy safety, fostering innovation that prioritizes user well-being and minimizes unintended negative consequences.
Appendix A Risk Taxonomy

### v2

@@
-A Risk Taxonomy for Evaluating AI-Powered Psychotherapy Agents
+A Risk Ontology for Evaluating AI-Powered Psychotherapy Virtual Agents
 Abstract.
@@
 Current evaluation techniques lack the sensitivity to detect subtle changes in patient cognition and behavior during therapy sessions that may lead to subsequent decompensation.
-We introduce a novel risk taxonomy specifically designed for the systematic evaluation of conversational AI psychotherapists.
-Developed through an iterative process including review of the psychotherapy risk literature, qualitative interviews with clinical and legal experts, and alignment with established clinical criteria (e.g., DSM-5) and existing assessment tools (e.g., NEQ, UE-ATR), the taxonomy aims to provide a structured approach to identifying and assessing user/patient harms.
-We provide a high-level overview of this taxonomy, detailing its grounding, and discuss potential use cases.
-We discuss two use cases in detail: monitoring cognitive model-based risk factors during a counseling conversation to detect unsafe deviations, in both human-AI counseling sessions and in automated benchmarking of AI psychotherapists with simulated patients.
-The proposed taxonomy offers a foundational step towards establishing safer and more responsible innovation in the domain of AI-driven mental health support.
-AI Psychotherapy, Risk Assessment, Intelligent Virtual Agents, LLM Safety, Mental Health Technology
-journalyear: 2025conference: ; May 2025; booktitle: Preprintccs: Applied computing Consumer healthccs: Information systems Language modelsccs: Human-centered computing Empirical studies in HCI
+We introduce a novel risk ontology specifically designed for the systematic evaluation of conversational AI psychotherapists.
+Developed through an iterative process including review of the psychotherapy risk literature, qualitative interviews with clinical and legal experts, and alignment with established clinical criteria (e.g., DSM-5) and existing assessment tools (e.g., NEQ, UE-ATR), the ontology aims to provide a structured approach to identifying and assessing user/patient harms.
+We provide a high-level overview of this ontology, detailing its grounding, and discuss potential use cases.
+We discuss four use cases in detail: monitoring real user interactions, evaluation with simulated patients, benchmarking and comparative analysis, and identifying unexpected outcomes.
+The proposed ontology offers a foundational step towards establishing safer and more responsible innovation in the domain of AI-driven mental health support.
+AI Psychotherapy, Risk Assessment, Intelligent Virtual Agents, LLM Safety, AI Alignment, Mental Health Technology
+journalyear: 2025conference: ; September 2025; booktitle: Preprintdoi: 10.1145/3717511.3749286isbn: 979-8-4007-1508-2/2025/09ccs: Applied computing Consumer healthccs: Information systems Language modelsccs: Human-centered computing Empirical studies in HCI
 1. Introduction
-Artificial Intelligence (AI) systems, particularly those leveraging Large Language Models (LLMs), are increasingly being developed and deployed as tools for mental health support, including entertainment chatbots ”masquerading” as therapists (Jurblum and Selzer, 2025; Barry, 2025; Jin et al., 2025; Na et al., 2025).
-The American Psychological Association has recently warned federal regulators about AI chatbots programmed to reinforce, rather than challenge, a user’s thinking, potentially driving vulnerable users to harm themselves or others (Lawrence et al., 2024; Barry, 2025).
-Reports have emerged, including instances where interactions with AI chatbot ”therapists” have been linked to severe negative outcomes.
-Notably, one case involved a 14-year-old boy’s suicide after months of interaction with chatbots on Character.AI, including a chatbot that falsely claimed to have several years’ experience as a licensed therapist (Roose, 2024; Barry, 2025).
-In this instance, the chabot reportedly reinforced the user’s harmful ideations rather than challenging them (Barry, 2025).
-Such events underscore the profound ethical responsibilities and potential harms associated with deploying these technologies to vulnerable populations, often with concerns raised about them being ”dangerous and untested” or lacking rigorous, context-appropriate safety evaluations (Roose, 2024; Barry, 2025; Stade et al., 2024b).
-A critical issue is the absence of standardized methods to evaluate these AI systems for the unique risks inherent in psychotherapeutic interactions.
-Existing LLM benchmarks often focus on general capabilities such as toxicity detection in isolated texts, failing to capture harms that may emerge from sustained, dynamic conversations (Chang et al., 2023; Zhang et al., 2025).
-Traditional clinical trials, while vital for efficacy, can be slow and ethically challenging for initial safety validation of unproven autonomous systems (Stade et al., 2024b; Heinz et al., 2025).
-This gap leaves users susceptible to unexpected adverse outcomes.
-To address this, we propose a risk taxonomy tailored for evaluating AI psychotherapists.
-We present a high-level overview of this taxonomy, derived from an iterative process including a review of the literature on psychotherapy risks, qualitative interviews with clinical and legal experts, and alignment with established clinical criteria and existing assessment tools.
-We hypothesize that this taxonomy can provide a structured and systematic approach to identifying, categorizing, and ultimately mitigating risks in AI-driven psychotherapy applications.
-Our primary aim is to provide a framework that can be used to evaluate automated psychotherapy systems for potential negative impacts before and during their deployment, fostering safer innovation in digital mental health.
+Intelligent Virtual Agents (IVAs), particularly those powered by Large Language Models (LLMs), are increasingly deployed as AI-powered psychotherapy agents, sometimes by entertainment chatbots ”masquerading” as qualified professionals (Jurblum and Selzer, 2025; Barry, 2025; Jin et al., 2025; Na et al., 2025).
+This trend raises serious safety concerns, highlighted by the American Psychological Association’s warning about AI reinforcing harmful thinking (Barry, 2025).
+Alarming incidents, such as a user’s suicide linked to interactions with an AI chatbot ”therapist” that reportedly amplified harmful ideations (Roose, 2024; Barry, 2025), underscore the profound risks.
+These events reveal the dangers of deploying such AI-powered IVAs, often ”dangerous and untested,” to vulnerable individuals without rigorous, context-appropriate safety evaluations (Roose, 2024; Barry, 2025; Stade et al., 2024).
+A critical challenge is the lack of standardized methods to evaluate these AI psychotherapy agents for risks unique to conversational therapeutic interactions.
+Current LLM benchmarks assess general capabilities like toxicity, but are ill-suited for the sustained, dynamic conversations inherent when AI-powered IVAs conduct therapy (Chang et al., 2024; Zhang et al., 2025).
+Traditional clinical trials, vital for efficacy, are also slow and ethically challenging for initial safety validation of autonomous AI psychotherapists (Stade et al., 2024; Heinz et al., 2025).
+This evaluation gap leaves users vulnerable to unforeseen adverse outcomes.
+To address this, we introduce a risk ontology specifically designed for evaluating AI psychotherapy agents.
+This paper presents a high-level overview of the ontology, developed through an iterative process involving psychotherapy risk literature review, qualitative interviews with clinical and legal experts, and alignment with established clinical criteria (e.g., DSM-5) and assessment tools (e.g., NEQ, UE-ATR).
+We posit that this ontology offers a structured, systematic approach to identify, categorize, and ultimately help to mitigate risks when IVAs deliver psychotherapy.
+Our primary aim is to provide a framework for evaluating these automated systems, fostering safer innovation in AI-driven mental health support.
 2. Related Work
@@
 (Mejía-Castrejón et al., 2024) adapted a medical model, classifying AEs by severity, relatedness, seriousness, and expectedness.
-Common AEs identified across studies and captured in assessment tools like the Negative Effects Questionnaire (NEQ) (Rozental et al., 2019) or the Unwanted Event to Adverse Treatment Reaction (UE-ATR) checklist (Linden, 2013) include negative emotions (e.g., anxiety, sadness), symptom worsening, the emergence of unpleasant memories, relationship strains, dependency on therapy, or reduced self-efficacy (Linden and Schermuly-Haupt, 2014; Yao et al., 2020).
+Common AEs identified across studies and captured in assessment tools like the Negative Effects Questionnaire (NEQ) (Rozental et al., 2016) or the Unwanted Event to Adverse Treatment Reaction (UE-ATR) checklist (Linden, 2013) include negative emotions (e.g., anxiety, sadness), symptom worsening, the emergence of unpleasant memories, relationship strains, dependency on therapy, or reduced self-efficacy (Linden and Schermuly-Haupt, 2014; Yao et al., 2020).
 Factors influencing these include patient and therapist characteristics, and specific techniques (Linden and Schermuly-Haupt, 2014).
@@
 Reviews show variability and inadequate reporting of negative effects in randomized clinical trials (Jonsson et al., 2014; Klatte et al., 2023), underscoring the need for standardized assessment.
-Our taxonomy development aims to build upon these existing efforts by providing a structured framework specially adapted for evaluating AI-powered psychotherapy agents.
+Our ontology development aims to build upon these existing efforts by providing a structured framework specifically adapted for evaluating AI psychotherapy agents.
 2.2. Safety Evaluation of LLMs in Healthcare
-The application of LLMs in high-stakes domains like healthcare, including psychotherapy, demands rigorous safety assessment due to the potential for significant harm (Stade et al., 2024b; Lawrence et al., 2024; Steenstra et al., 2024).
-General LLM safety evaluations often focus on risks including bias, toxicity, or misinformation (Weidinger et al., 2021; Ji et al., 2023), using benchmarks such as ALERT (Tedeschi et al., 2024) or SafetyBench (Zhang et al., 2023).
+The application of LLMs in high-stakes domains like healthcare, including psychotherapy, demands rigorous safety assessment due to the potential for significant harm (Stade et al., 2024; Lawrence et al., 2024; Steenstra et al., 2024).
+General LLM safety evaluations often focus on risks including bias, toxicity, or misinformation (Weidinger et al., 2022; Ji et al., 2025), using benchmarks such as ALERT (Tedeschi et al., 2024) or SafetyBench (Zhang et al., 2024).
 However, these may not adequately capture the interactive and relational risks specific to psychotherapy.
-Safety frameworks are emerging for AI in mental health, including the READI framework (Stade et al., 2024a) and the BPES checklist, emphasizing ”Do No Harm” (Khan and Seto, 2023).
+Safety frameworks are emerging for AI in mental health, including the READI framework (Stade et al., 2025) and the BPES checklist, emphasizing ”Do No Harm” (Khan and Seto, 2023).
+Complementing these expert-driven approaches, some research grounds risk assessment in the direct experiences of users.
+For instance, Chandra et al.
+(Chandra et al., 2025) developed a psychological risk taxonomy by surveying 283 individuals with lived mental health experience about their negative interactions with AI conversational agents, categorizing risks into specific AI behaviors, negative psychological impacts, and the user contexts that mediate them.
 Dialogue safety in mental health support is also a focus (Qiu et al., 2023).
 For instance, $\Psi$-ARENA offers an interactive platform for evaluating LLM-based psychological counselors (Zhu et al., 2025).
-Despite these advances, a standardized taxonomy to systematically identify and categorize the spectrum of potential risks specific to AI psychotherapeutic interactions, grounded in clinical realities and existing harm assessment principles, is still lacking.
+Despite these advances, a standardized ontology is still lacking to operationalize risk detection by systematically assessing and categorizing the spectrum of potential harms specific to AI psychotherapeutic interactions, grounded in clinical realities and existing assessment principles.
 This is crucial as AI systems, unlike human therapists, may lack nuanced understanding of human emotion, contextual subtleties, and the ability to dynamically repair relational ruptures, potentially leading to unique failure modes.
-3. Taxonomy Development
-The development of our risk taxonomy followed an iterative design process.
-We first formulated a preliminary taxonomy based on an in-depth literature review of psychotherapy risk (e.g., (Linden, 2013; Rozental et al., 2018; Klatte et al., 2023)).
+3. Ontology Development
+The development of our risk ontology followed an iterative design process.
+We first formulated a preliminary ontology based on an in-depth literature review of psychotherapy risk (e.g., (Linden, 2013; Rozental et al., 2018, 2016; Klatte et al., 2023)).
 This initial framework subsequently underwent expert review through semi-structured interviews with 11 experts: 10 practicing psychologist professionals and one legal professional specializing in healthcare malpractice.
-These interviews were designed to explore perceptions of risk in human therapy, identify potential failure modes for AI therapists, and gather direct feedback on the preliminary taxonomy.
+These interviews were designed to explore perceptions of risk in human therapy, identify potential failure modes for AI therapists, and gather direct feedback on the preliminary ontology.
 The interview study was approved by Northeastern University’s IRB.
-The interviews were analyzed using thematic analysis (Clarke and Braun, 2017) to reveal six key themes.
-3.1. Theme 1: Therapy may require short-term discomfort
-The first theme highlighted the difficulty in defining and identifying ”harm” within the therapeutic context itself.
+The interviews were analyzed using thematic analysis (Clarke and Braun, 2017) to reveal six key themes that directly informed the final structure and content of the ontology.
+The interview study was approved by Northeastern University’s IRB.
+Theme 1: Therapy may require short-term discomfort The first theme highlighted the difficulty in defining and identifying ”harm” within the therapeutic context itself.
 Unlike the clearer ”do no harm” principle in general medicine, psychotherapy often involves navigating ”intentional discomfort” as part of the therapeutic process.
@@
 Furthermore, harm can be subjective, context-dependent (”contextual based on your history”), and may not manifest ”in the moment,” only emerging later through ”rumination.”
-3.2. Theme 2: Therapist risk factors
-Therapist factors related to risk are significant, encompassing competence (”evidence-based care,” ”qualifications,” ”knowing their specialty”), ethical adherence (”violates boundaries,” ”confidentiality”), and skillful technique (avoiding ”leading questions,” ”invalidating” responses, ”negative vocabulary,” ”overly reassuring”).
+Theme 2: Therapist risk factors Therapist factors related to risk are significant, encompassing competence (”evidence-based care,” ”qualifications,” ”knowing their specialty”), ethical adherence (”violates boundaries,” ”confidentiality”), and skillful technique (avoiding ”leading questions,” ”invalidating” responses, ”negative vocabulary,” ”overly reassuring”).
 Therapist ”biases,” ”personal opinions,” lack of ”cultural awareness,” ”distractions,” and ”burnout” were cited as potential sources of unintentional harm.
 Building ”therapeutic rapport” and ”trust” was seen as a primary mitigator, allowing for ”repair” if harm occurs.
-3.3. Theme 3: Patient risk factors
-Patient vulnerability and context also play a critical role.
+Theme 3: Patient risk factors Patient vulnerability and context also play a critical role.
 ”Individual differences” mean risk factors vary; internal states (”hopelessness,” ”anxiety”), patient history (”memories,” ”interpretations”), the ”social environment,” ”physical environment” (including ”access to means”), and ”protective factors” all interact dynamically.
-3.4. Theme 4: Real-time nature of risk assessment
-Identifying risks in real-time relies heavily on a therapist’s ability for ”moment-to-moment tracking,” interpreting subtle ”nonverbal body language,” ”facial expressions,” and ”para verbal cues” (e.g., ”intonation”), alongside verbal content and formal ”assessment tools” (e.g., ”Columbia Scale”(Posner et al., 2008), ”PHQ-9”(Kroenke et al., 2001)).
+Theme 4: Real-time nature of risk assessment Identifying risks in real-time relies heavily on a therapist’s ability for ”moment-to-moment tracking,” interpreting subtle ”nonverbal body language,” ”facial expressions,” and ”para verbal cues” (e.g., ”intonation”), alongside verbal content and formal ”assessment tools” (e.g., ”Columbia Scale”(Posner et al., 2008), ”PHQ-9”(Kroenke et al., 2001)).
 However, assessment is imperfect due to patient ”concealment,” the limits of prediction (”intention could change”), and the significant information loss in modalities lacking visual and auditory cues, such as ”texting” therapy, which was viewed as particularly risky by many participants.
 Critically, effective responses require understanding the ”underlying cause,” not just the presenting ”symptom.”
-3.5. Theme 5: Perceived shortcomings of AI therapists
-When considering AI therapists, participants largely believed the fundamental risks in human therapy remain the same but are potentially ”higher” or exacerbated.
+Theme 5: Perceived shortcomings of AI Psychotherapists When considering AI therapists, participants largely believed the fundamental risks in human therapy remain the same but are potentially ”higher” or exacerbated.
 The primary concern stems from the AI’s inherent inability (especially in text-based interactions) to perceive the rich nonverbal and paralinguistic data humans use to gauge internal states, build rapport, and understand nuance (”missing the person-to-person component,” ”can’t synthesize,” loss of ”nuance”).
@@
 Interestingly, AI was often considered potentially ”better” than human texting therapy, as the latter shares AI’s lack of nonverbals without AI’s potential processing advantages.
-3.6. Theme 6: Refinement of risk taxonomy
-The feedback on the proposed risk taxonomy indicated it was a ”good start” covering many ”big risks,” but highlighted needs for refinement.
-Key omissions noted were ”trauma” and ”abuse.” Participants stressed the necessity of integrating ”environmental” and ”social factors,” the ”timeline” of symptoms, and grounding factors in established tools like the ”DSM” or clinical ”inventories,” while acknowledging the limitations of categorical approaches versus ”individual differences” and ”nuance.” The taxonomy’s current structure was seen as potentially useful but insufficient alone for risk assessment without incorporating deeper context and validated measures.
-These insights drove significant revisions: patient factors within the risk taxonomy were linked to the DSM-5, factor values are now compared to baselines to accommodate individual differences, and risk categorization was grounded in approaches similar to the NEQ and UE-ATR.
-4. Overview of the Risk Taxonomy
-The refined risk taxonomy, derived from literature (Linden, 2013; Rozental et al., 2018; Klatte et al., 2023), expert interviews, and DSM-5 criteria (Black and Grant, 2014), provides a structured framework for identifying potential negative outcomes in AI psychotherapy.
-The full taxonomy can be found in Appendix A.
-It has two main categories:
+Theme 6: Refinement of risk ontology The feedback on the proposed risk ontology indicated it was a ”good start” covering many ”big risks,” but highlighted needs for refinement.
+Key omissions noted were ”trauma” and ”abuse.” Participants stressed the necessity of integrating ”environmental” and ”social factors,” the ”timeline” of symptoms, and grounding factors in established tools like the ”DSM” or clinical ”inventories,” while acknowledging the limitations of categorical approaches versus ”individual differences” and ”nuance.” The ontology’s current structure was seen as potentially useful but insufficient alone for risk assessment without incorporating deeper context and validated measures.
+These insights drove significant revisions to the ontology’s structure and granularity.
+Specifically, the theme that ”Therapy may require short-term discomfort” led to the crucial distinction in our ontology between ”In-Session Warning Signs” (which may include therapeutically necessary negative affect) and ”Potential Real-World Consequences” (clear harms).
+This ensures the evaluation framework does not penalize an AI for inducing transient discomfort that is part of a valid therapeutic process.
+Furthermore, the themes of ”Patient risk factors” and the need to account for ”individual differences” directly informed our decision to incorporate baseline comparisons for psychological constructs, allowing for a more personalized risk assessment.
+Finally, feedback on the need to ground the ontology in established tools led to the explicit alignment of our ”In-Session Warning Signs” and ”Potential Real-World Consequences”, with concepts from the DSM-5 and existing instruments like the NEQ and UE-ATR.
+4. Overview of the Risk Ontology
+The refined risk ontology (see Appendix A), derived from literature (e.g., (Linden, 2013; Rozental et al., 2018, 2016; Klatte et al., 2023)), expert interviews, and DSM-5 criteria (Association et al., 2013), offers a structured framework to identify potential negative outcomes in AI psychotherapy.
+It has three overarching categories:
 1.
-Immediate Risk: This category encompasses situations of imminent danger to the user or others, requiring urgent intervention.
-Examples include explicit suicidal intent with plan and means, threats of harm to others, or severe psychological decompensation such as acute psychosis.
-Identification relies primarily on dialogue-based triggers.
-Evaluation assesses the AI’s ability to recognize these triggers and respond according to predefined, ethically-informed action plans, such as attempting de-escalation, recommending emergency resources (e.g., 911, 988 Lifeline), and contacting/flagging a human supervisor immediately.
+Acute Crisis: This category represents immediate and severe danger to the user that requires urgent intervention.
+Identification is entirely dialogue-based, evaluating the AI psychotherapists ability to recognize and respond to scenarios such as clear suicidal intent, threats of harm to others, or severe psychological decompensation.
+The framework assesses the AI’s performance against a predefined four-step action plan: (1) Assess, (2) De-escalate, (3) Recommend Emergency Services, and (4) Request Human Consultation.
 2.
-Potential Risk: This category refers to emerging states or interaction patterns suggesting heightened vulnerability to negative therapeutic outcomes, functional deterioration, or adverse impacts on the therapeutic process, often aligning with unwanted events in the NEQ or UE-ATR (Rozental et al., 2019; Linden, 2013).
-Key subcategories are: 2.1.
-Symptom Exacerbation, Emergence of New Symptoms, or Maladaptive Coping, which involves worsening of clinical symptoms (e.g., increased negative thoughts or feelings) or new maladaptive behaviors (e.g., Worry/Rumination Frequency); 2.2.
-Patient Destabilization or Feeling Overwhelmed, indicating the user is struggling with the therapeutic process (e.g., increased confusion, emotional lability); 2.3.
-Therapeutic Alliance Damage or Negative Relational Dynamics, covering negative shifts in the user-AI relationship including mistrust, feeling judged, or over-dependence; and 2.4.
-Reduced Engagement or Potential for Premature Termination, signaling decreased participation or commitment within the therapeutic process or current therpay session.
-Identification of these potential risks involves tracking dynamic changes in internal cognitive, emotional, and behavioral factors, which are conceptualized as scales (e.g., from 1-low to 10-high for intensity or frequency, such as ”Hopelessness Level” or ”Collaboration/Engagement Level”).
-The up (↑) and down (↓) arrows accompanying these factors in the full taxonomy (see Appendix A) specify the direction of change from a user’s baseline state that may signify an increased potential risk.
-For example, an increase (↑) in ”Hopelessness Level” or a decrease (↓) in ”Collaboration/Engagement Level” would be flagged as potentially problematic.
-Evaluation should assess harmful shifts in these factors from baseline values, allowing for the tracking of subtle deteriorations or emerging vulnerabilities.
-While empirically-grounded numeric thresholds for what constitutes a significant change is a goal for future work, the current approach suggests that any major deviation beyond simple, minor fluctuations around a baseline cognitive factor value (e.g., more than +/- 1 point on a 1-10 scale from the baseline), or other abnormal fluctuations, could indicate an increased potential risk requiring attention.
-Furthermore, as expert participants highlighted, these indicators should ideally be monitored over the course of a session or across multiple interactions to understand their trajectory and contextual significance, rather than relying solely on isolated measurements.
-For instance, a consistent decrease (↓) in the Motivation/Self-Efficacy Level factor over multiple sessions could strongly suggest an increased risk for Reduced Engagement / Potential for Premature Termination.
-5. Expected Use Cases and Adaptability
-The taxonomy standardizes safety evaluation criteria for AI-powered psychotherapy agents. There are several ways the taxonomy can be used to improve safety. We outline several use cases here.
-5.1. Monitoring Real User Interactions
-The risk taxonomy could be used to identify emergent risks in real-time, as a user interacts with an AI therapist, for example to provide early termination of a harmful session, or feedback to developers or regulatory bodies.
-Monitoring should be performed with ethical consent, and could be accomplished by tracking user risk factors from pre- to post-interaction to identify ”Potential Risk” shifts, while dialogue tools scan for ”Immediate Risk” triggers.
-5.2. Evaluation with Simulated Patients
-The risk taxonomy could also be used to ”monitor” interactions between an AI therapist and a simulated patient, such as (Steenstra et al., 2025), providing a powerful evaluation testbed that can be used to assess systems for safety and potential harm.
-During these simulated interactions, simulated patient agents’ dynamic internal states corresponding to ”Potential Risk” factors (e.g., ”Hopelessness Level) are updated.
-This allows systematic testing by comparing baseline to post-interaction values of internal patient factors and observing AI responses to diverse personas, a methodology previously validated in our development of SimPatient (Steenstra et al., 2025), an LLM-powered system where embodied simulated patients featured dynamic internal cognitive states (e.g., self-efficacy) that evolved during interactions.
-5.3. Benchmarking and Comparative Analysis
-The risk taxonomy provides a standardized framework essential for robust benchmarking and comparative analysis of different AI psychotherapist models or system versions.
-By systematically evaluating AI therapists against the defined risk categories and factors, consistent and comparable risk profiles can be generated for each system.
-This process can highlight the relative strengths and weaknesses of various AI approaches in terms of their propensity to elicit or mitigate specific types of risks.
-For example, one model might excel at de-escalating immediate suicidal ideation but perform poorly by increasing therapeutic alliance damage, while another might show the opposite pattern.
-Such comparative insights are invaluable for developers seeking to enhance model safety, for organizations making informed deployment decisions based on specific use-case risk tolerances, and for fostering a more transparent and accountable development ecosystem focused on minimizing harm.
-5.4. Identifying Unexpected Outcomes
-Finally, the risk taxonomy serves as a crucial tool for identifying unexpected or emergent adverse outcomes that may arise during automated therapy.
-This is achieved through systematic monitoring of user interactions and therapeutic outcomes, continually assessed against the taxonomy’s defined risk factors and categories.
-Such monitoring can reveal significant deviations from established baseline values—which ideally exhibit only safe fluctuations—or flag the emergence of previously unrecognized harmful patterns.
-When substantial deviations or novel harmful patterns are detected, a deeper analysis is warranted.
-This involves scrutinizing the conversational context surrounding these occurrences to understand the specific interaction dynamics or AI behaviors that may have precipitated them.
-Uncovering these root causes is invaluable, as it directly informs targeted safety improvements and iterative refinements of the AI model, ultimately leading to a more robust and safer therapeutic tool.
-While our risk taxonomy provides a general framework, it can also be augmented for specific populations or problems.
-For instance, evaluating an AI psychotherapist for helping patients with substance use disorders could add ”Potential Risk” factors such as ”Substance Craving Intensity”, ”Ambivalence about changing substance use” ”Substance Use Frequency/Amount,” ”Level of Intoxication/Withdrawal Symptoms”, and relational dynamics like ”Dishonesty/Minimization regarding substance use.” This tailoring allows for more nuanced risk assessment.
-Future work includes validating the taxonomy against human expert judgments.
+In-Session Warning Signs: This category tracks dynamic, moment-to-moment changes in the patient’s internal psychological state that may predict future harm or signal heightened vulnerability.
+These are not direct harms but are leading indicators.
+The framework operationalizes key psychological constructs (e.g., Hopelessness, Self-Efficacy) as internal state variables, which are conceptualized as a 5-point Likert scale.
+These constructs, are dynamically updated in response to the AI psychotherapist’s dialogue.
+Evaluation focuses on tracking the magnitude and direction of these changes, acknowledging that while some transient negative shifts can be part of the therapeutic process.
+3.
+Potential Real-World Consequences: This category represents the tangible, adverse outcomes that could manifest in the patient’s life following a therapeutic interaction, often aligning with NEQ and UE-ATR unwanted events (Rozental et al., 2016; Linden, 2013).
+Unlike the internal states of ”In-Session Warning Signs”, these are the observable ”harms” or deteriorations that serve as the primary dependent variables for evaluating the AI’s long-term safety and efficacy.
+Each potential consequence is causally linked to changes in the ”In-Session Warning Signs”.
+The ontology aims to standardize safety evaluation for AI psychotherapists. We outline several use cases for improving safety:
+Monitoring Real User Interactions: The ontology can identify emergent risks in real-time during user-AI psychotherapist interactions, enabling early termination of harmful sessions or providing feedback to developers and regulators.
+Monitoring, with ethical consent, could track a user’s psychological construct intensities pre- to post-interaction for ”In-Session Warning Signs” shifts, while dialogue tools scan for ”Acute Crisis” events.
+Evaluation with Simulated Patients: The risk ontology can also structure the monitoring of interactions between an AI psychotherapist and a simulated patient.
+This approach is exemplified in our work on SimPatient, a testbed designed for safety and quality of care assessment (Steenstra et al., 2025).
+Within this system, the simulated patient is equipped with dynamic internal states (e.g., Self-Efficacy) that directly correspond to the ontology’s ”In-Session Warning Signs”.
+This design enables systematic testing by comparing the patient’s internal states throughout the interaction, allowing for quantitative assessment of the AI psychotherapist’s impact on diverse personas.
+Benchmarking and Comparative Analysis: The ontology provides a standardized framework for robust benchmarking and comparing AI psychotherapist models and versions.
+Systematically evaluating AI psychotherapists against risk categories generates consistent, comparable risk profiles, highlighting strengths and weaknesses in eliciting and mitigating specific risks.
+Such insights could aid developers in enhancing safety and fostering a transparent development ecosystem focused on minimizing harm.
+Identifying Unexpected Outcomes: The ontology is crucial for identifying unexpected and emergent adverse outcomes in automated therapy.
+Systematic monitoring of user interactions and outcomes against ontology categories can reveal significant deviations from baselines or flag harmful patterns.
+Substantial deviations warrant deeper analysis, scrutinizing conversational context to understand precipitating AI behaviors.
+Uncovering root causes informs targeted safety improvements and AI model refinements.
+Our generalized risk ontology can also be augmented for specific populations or problems.
+For instance, evaluating an AI psychotherapist for patients with substance use disorders, could add a ”In-Session Warning Signs” factor like ”Substance Craving Intensity”.
+This tailoring enables nuanced risk assessment.
+A critical next step is the formal validation of the ontology itself.
+We envision a multi-pronged approach involving content validation by clinical experts, construct validation by correlating its categories with validated psychometric surveys, and empirical validation by establishing inter-rater reliability against human-rated therapeutic dialogues.
 5. Conclusion
@@
 Documented negative outcomes demand standardized evaluations for unique interaction risks.
-We developed the risk taxonomy, from an iterative process involving literature, expert insights, and DSM-5, and harm checklists.
+We developed the risk ontology, from an iterative process involving literature, expert insights, and DSM-5, and harm checklists.
 It offers a structured approach to identifying and categorizing harms.
-Its overview, grounding, and adaptable use cases (tracking risk factor changes from baseline to post-interaction) aim to equip stakeholders to ensure AI psychotherapy safety, fostering innovation that prioritizes user well-being and minimizes unintended negative consequences.
-Appendix A Risk Taxonomy
+Its overview, grounding, and adaptable use cases (tracking in-session warning signs) aim to equip stakeholders to ensure AI psychotherapy safety, fostering innovation that prioritizes user well-being and minimizes unintended negative consequences.
+Appendix A Supplementary Material – AI Therapy Risk Ontology
+Figure 1. A High-Level Overview of the AI Therapy Risk Ontology:
+Figure 2. Action Plans for Responding to Acute Crisis Scenarios:
+Figure 3. Definitions of Psychological Constructs Used as In-Session Warning Signs:
+Figure 4. Potential Behavioral Harms and Their Predictive Psychological Constructs:
+Figure 5. Potential Cognitive and Affective Harms and Their Predictive Psychological Constructs:
+Figure 6. Potential Relational Harm and its Predictive Psychological Constructs: